Calculate the moment of inertia of the composite area below about: a) The x- and y-axes shown (Ix and ly). b) Thex- and y-axePlease show all work and formulas used to solve. Thank you for your help.
Answer #1

160 140 mm 1 2 200 40 3 40 mm

a) The Moment of Area about x axis for Area 1.

-40! * 40 +120 * 40 * (200-20). 1.5616 * 108mm 12

The Moment of Area about x axis for Area 2.

40 200 12 +20040 1002 1.067108mm 1r2

The Moment of Area about x axis for Area 3.

(120 40)405 12 80 * 40 * 202 1.067 * 10 mrn4

Total Moment of Area about x Axis

Ix= Ix1+Ix2+Ix3 = 2.735*108 mm4

The Moment of Area about y axis for Area 1.

(160 - 40) 40 12 +12040 (40 +120/2)2 5.376 10mm1

The Moment of Area about x axis for Area 2.

200 403 12 200 40204.267 10mm y2

The Moment of Area about x axis for Area 3.

(120 - 40) 40 12 +80 40 (4080/2)2 2.219 10mm

Total Moment of Area about x Axis

Iy= Iy1+Iy2+Iy3 = 8.0214*107 mm4

b) Centroid of Area 1 (x1,y1) (120,180)

Centroid of Area 2 (x2,y2) (20,100)

Centroid of Area 3 (x3,y3) (80,20)

T1A1 2A2 .T3A3120 * 120 40+ 20 200 40 80 8040 40 12020040+80 40

\bar{x} = 62

yi Ai y2A2 УзАз 180 * 120% 40 + 100 +200% 40 + 20 * 80 * 40 40 120200 4080 40

\bar{y} = 108

li = Irt(A1+A2+Аз )ゾー2.735 * 108 + (40% 120 +200+40+80 * 40) * 1082

I 4.60124 108mm4

I_{\bar{y}} = I_y +(A_1+A_2+A_3)\bar{x}^2=8.0214*10^7 +(40*120+200*40+80*40)*62^2

I_{\bar{y}} = 1.4172*10^8 mm^4